Answer the following simple interest questions:

1. Jeff borrowed $25 from his dad to buy a fishing license and promised to pay his dad back in one month. His dad agreed, but said that Jeff must pay 15% annual interest. How much interest will Jeff pay?

2. Anna Maria applied for and received a loan to pay for college. She borrowed $25,000. She must pay the loan back in five years with annual interest of 4%. How much money will she pay when she pays off the loan?

3. Daunell purchased a new TV on September 15th. The cost of the TV was $2,750 and she has until January 15th to pay that amount to the retailer. The retailer has added 21% annual interest to the price of the TV. When Daunell pays off the TV on January 15th, how much will she pay?

4. Joe bought a boat from his brother on March 1st with a promise to write a check for the price of the boat, $5,000, on March 4th. On March 2nd Joe lost his job so his brother said he could pay the money when he found a new job. Six months later Joe still didn’t have a job, but he was tan from a lot of water skiing. Joe’s brother said that Joe needed to pay for the boat in 3 months at 5% annual interest. How much will Joe pay his brother in 3 months?

5. Mary borrowed $5 from her older brother and paid him back $7 dollars three days later. What annual interest rate did her older brother charge her?

6. Rachel, Mary’s sister, also borrowed money from her older brother. She repaid him $10 three days after he loaned her money. He charged her 75% annual interest. What amount did Rachel borrow?
Answer #1
Simple Interest =Principal*Interest Rate*Time in year
1 Principal $                                           25.00
Interest Rate 15%
Time 1 month
Interest $                                             0.31 =25*15%*1/12
2 Principal $                                  25,000.00
Interest Rate 4%
Time                                                 5.00 years
Interest $                                     5,000.00 =25000*4%*5
Money to be paid (Principal +interest) $                                  30,000.00 =25000+5000
3 Principal $                                     2,750.00
Interest Rate 21%
Time                                             123.00 days
Interest $                                        194.61 =2750*21%*123/365
Money to be paid (Principal +interest) $                                     2,944.61 =2750+194.61
4 Principal 5000
Interest Rate 5%
Time 9 months
Interest $                                        187.50 =5000*5%*9/12
Money to be paid (Principal +interest) $                                     5,187.50 =5000+187.5